The length of a rectangle is 3 centimeters more than three times the width. If the perimeter of the rectangle is 22 centimeters, what are its dimensions?
Found 2 solutions by kelsdf, Alan3354:
Answer by kelsdf(2)

Let's make the width equal x.
That would make the length 3x+3.
If the perimeter is 22, the formula is length+length+width+width or 2length+2width.
So we would plug in our variables giving us the equation: 2x+[2(3x+3)]=22.
Multiply and you'd get 2x+6x+6=22.
Subtract 4 from each side to get 2x+6x=16.
Add like terms giving you: 8x=16
Divide both sides 8 to get x=2.
Plug in 6 to get x=6 and 3x+3=21.
So your length would be 21 and your width would be 6.

Answer by Alan3354(31532)

The length of a rectangle is 3 centimeters more than three times the width. If the perimeter of the rectangle is 22 centimeters, what are its dimensions?
------------
L = 3W + 3
2L + 2W = 22
Sub for L into eqn 2
2(3w+3) + 2W = 22
3W+3 + W = 11
4W = 8
W = 2 cm
-----------
L = 9 cm